# Addressing Loss of Plasticity and Catas- trophic Forgetting in Continual Learning

### v1

@@ 3 Method
To define utility precisely, let us consider that the learner produces the predicted output $\hat{{\bm{y}}}$ using a neural network with $L$ layers, parametrized by the set of weights $\mathcal{W}=\{{\bm{W}}_{1},...,{\bm{W}}_{L}\}$. Here ${\bm{W}}_{l}$ is the weight matrix at the $l$-th layer, and its element at the $i$-th row and the $j$-th column is denoted by $W_{l,i,j}$. At each layer $l$, we get the activation output ${\bm{h}}_{l}$ of the features by applying the activation function $\bm{\sigma}$ to the activation input ${\bm{a}}_{l}$: ${\bm{h}}_{l}=\bm{\sigma}({\bm{a}}_{l})$. We simplify notations by defining ${\bm{h}}_{0}\doteq{\bm{x}}$. The activation output ${\bm{h}}_{l}$ is then multiplied by the weight matrix ${\bm{W}}_{l+1}$ of layer $l+1$ to produce the next activation input: ${a}_{l+1,i}=\sum_{j=1}^{d_{l}}{{W}_{l+1,i,j}}{h}_{l,j},\forall i$, where ${\bm{h}}_{l}\in\mathbb{R}^{d_{l}}$. Here, $\bm{\sigma}$ applies activation element-wise for all layers except for the final layer, which becomes the softmax function.
@@ 3.2 Utility-based Perturbed Gradient Descent (UPGD)
Now, we devise a new method called Utility-based Perturbed Gradient Descent (UPGD) that performs gradient-based learning guided by utility-based information. The utility information is used as a gate, referred to as utility gating, for the gradients to prevent large updates to already useful weights, addressing forgetting. On the other hand, the utility information helps maintain plasticity by perturbing unuseful weights which become difficult to change through gradients (see Dohare et al. 2023a). The update rule of UPGD is given by
@@
where $\xi\sim\mathcal{N}(0,1)$ is noise, $\alpha$ is the step-size parameter, and $\bar{U}_{l,i,j}\in[0,1]$ is a scaled utility. For important weights with utility $\bar{U}_{l,i,j}=1$, the weight remains unaltered even by gradient descent, whereas unimportant weights with $\bar{U}_{l,i,j}=0$ get updated by both perturbation and gradient descent.
@@ 4.2 UPGD Against Loss of Plasticity
We compare SGDW, PGD, S&P, which addresses loss of plasticity, Adam with weight decay (Loshchilov & Hutter 2019) known as AdamW, UPGD-W, and Non-protecting UPGD-W. We also introduce and compare against Streaming Elastic Weight Consolidation (S-EWC), Streaming Synaptic Intelligence (S-SI), and Streaming Memory-Aware Intelligence (S-MAS). These methods can be viewed as a natural extension of EWC (Kirkpatrick et al. 2017), SI (Zenke et al. 2017), and MAS (Aljundi et al. 2018), respectively, which are regularization-based methods for mitigating forgetting to the streaming learning setting. Finally, we introduce and compare against Streaming RWalk (S-RWalk). This can be seen as a natural extension of RWalk (Chaudhry et al. 2018), a method that addresses both issues, adapted for streaming learning. We write the update rule of the last four methods in streaming learning as $w_{l,i,j}\leftarrow w_{l,i,j}-\alpha\left[\nicefrac{{\partial\mathcal{L}}}{{\partial w_{l,i,j}}}+\kappa\Omega_{l,i,j}(w_{l,i,j}-\bar{w}_{l,i,j})\right]$, where $\kappa$ is the regularization factor, $\Omega_{l,i,j}$ is the weight importance, and $\bar{w}_{l,i,j}$ is the trace of weight $i,j$ at the $l$-th layer. The weight importance is estimated as a trace of squared gradients in S-EWC, whereas it is estimated as a trace of gradient magnitudes in S-MAS. S-RWalk is different from RWalk since it uses a trace of previous weights instead of the instantaneous previous weights. We omit the details of S-SI and RWalk weight importance estimation here and defer it to Appendix I.4. Lastly, since we compare against methods that use first-order information, we use global first-order utility traces in this and subsequent experiments for a fair comparison (however, see Appendix H for an experiment using second-order utilities).
Refer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix A Convergence Analysis for UPGD and Non-protecting UPGD
We further assume that $\mathcal{L}$ has bounded variance in the gradients $\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta},S)-\nabla f(\bm{\theta})\|^{2}]\leq\sigma^{2},\forall\bm{\theta}\in\mathbb{R}^{d}$. Note that the assumption of L-smoothness on the sample loss result in L-smooth expected loss too, which is given by $\|\nabla f(\bm{\theta}_{1})-\nabla f(\bm{\theta}_{2})\|\leq L\|\bm{\theta}_{1}-\bm{\theta}_{2}\|$. We assume that the perturbation noise in both UPGD versions has bounded variance $\mathbb{E}[\|\bm{\xi}\|^{2}]\leq\sigma^{2}_{n}$. For the simplicity of this proof, we use the true instantaneous weight utility, not an approximated one. We assume that the utility of any connection in the network is upper bounded by a number close to one $\bar{u}\to 1^{-}$.
@@ A.1 Non-protecting Utility-based Perturbed Gradient Descent
$\mathbb{E}[\|{\bm{g}}_{t}\|^{2}]$ $\leq\frac{\sigma^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2},$
@@
$\mathbb{E}_{t}[f(\bm{\theta}_{t+1})|\bm{\theta}_{t}]$ $\leq f(\bm{\theta}_{t})-\alpha\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2}\right)$
$=f(\bm{\theta}_{t})-\frac{2\alpha-L\alpha^{2}}{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}\right).$
@@
$\frac{2\alpha-L\alpha^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq f(\bm{\theta}_{1})-\mathbb{E}[f(\bm{\theta}_{T+1})]+\frac{L\alpha^{2}T}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}\right).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq 2\frac{f(\bm{\theta}_{1})-f(\bm{\theta}^{*})}{T(2\alpha-L\alpha^{2})}+\frac{L\alpha^{2}2(\sigma^{2}_{n}b_{t}+\sigma^{2})}{b_{t}(2\alpha-L\alpha^{2})}.$
@@ A.2 Utility-based Perturbed Gradient Descent
$\leq f(\bm{\theta}_{t})-\alpha\bar{\rho}\sum_{i=1}^{d}\nabla[f(\bm{\theta}_{t})]_{i}^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2}\right)$
$=f(\bm{\theta}_{t})-\left(\alpha\bar{\rho}-\frac{L\alpha^{2}}{2}\right)\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}\right).$
@@
$\frac{2\alpha\bar{\rho}-L\alpha^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq f(\bm{\theta}_{1})-\mathbb{E}[f(\bm{\theta}_{T+1})]+\frac{L\alpha^{2}T}{2}\left(\sigma^{2}_{n}+\frac{\sigma^{2}}{b_{t}}\right).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq 2\frac{f(\bm{\theta}_{1})-f(\bm{\theta}^{*})}{T(2\alpha\bar{\rho}-L\alpha^{2})}+\frac{L\alpha^{2}(\sigma^{2}_{n}b_{t}+\sigma^{2})}{b_{t}(2\alpha\bar{\rho}-L\alpha^{2})}.$
@@ Appendix B Utility Propagation
$f_{l,i,j}$ $\doteq\frac{\sigma^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}f_{l+1,k,i},$
$s_{l,i,j}$ $\doteq\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}2s_{l+1,k,i}\frac{\sigma^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}.$
@@
$\frac{\partial\mathcal{L}}{\partial a_{l,i}}$ $=\sum_{k=1}^{|{\bm{a}}_{l+1}|}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}\frac{\partial a_{l+1,k}}{\partial h_{l,i}}\frac{\partial h_{l,i}}{\partial a_{l,i}}=\sigma^{\prime}(a_{l,i})\sum_{k=1}^{|{\bm{a}}_{l+1}|}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i},$ (9)
@@
$\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l,i}}}}$ $\doteq\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{[}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma^{\prime\prime}(a_{l,i})\Bigg{]},$ (11)
@@
$f_{l,i,j}$ $=-\sigma^{\prime}(a_{l,i})h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\Bigg{)}$ (14)
$s_{l,i,j}$ $=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma^{\prime\prime}(a_{l,i})\Bigg{)}$ (15)
@@
$f_{l,ij}$ $=-\sigma^{\prime}(a_{l,i})h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\Bigg{)}$
$=\frac{\sigma^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}-\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}h_{l,i}W_{l+1,k,i}\Bigg{)}$
$=\frac{\sigma^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}f_{l+1,k,i},$ (16)
$s_{l,i,j}$ $=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma^{\prime\prime}(a_{l,i})\Bigg{)}$
$=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}h^{2}_{l,i}W^{2}_{l+1,k,i}\frac{\sigma^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}$
$=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}2s_{l+1,k,i}\frac{\sigma^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}.$ (17)

### v2

@@ 3 Method
To define utility precisely, let us consider that the learner produces the predicted output $\hat{{\bm{y}}}$ using a neural network with $L$ layers, parametrized by the set of weights $\mathcal{W}=\{{\bm{W}}_{1},...,{\bm{W}}_{L}\}$. Here ${\bm{W}}_{l}$ is the weight matrix at the $l$-th layer, and its element at the $i$-th row and the $j$-th column is denoted by $W_{l,i,j}$. At each layer $l$, we get the activation output ${\bm{h}}_{l}$ of the features by applying the activation function $\bm{\sigma}_{a}$ to the activation input ${\bm{a}}_{l}$: ${\bm{h}}_{l}=\bm{\sigma}_{a}({\bm{a}}_{l})$. We simplify notations by defining ${\bm{h}}_{0}\doteq{\bm{x}}$. The activation output ${\bm{h}}_{l}$ is then multiplied by the weight matrix ${\bm{W}}_{l+1}$ of layer $l+1$ to produce the next activation input: ${a}_{l+1,i}=\sum_{j=1}^{d_{l}}{{W}_{l+1,i,j}}{h}_{l,j},\forall i$, where ${\bm{h}}_{l}\in\mathbb{R}^{d_{l}}$. Here, $\bm{\sigma}_{a}$ applies activation element-wise for all layers except for the final layer, which becomes the softmax function.
@@ 3.2 Utility-based Perturbed Gradient Descent (UPGD)
Now, we devise a new method called Utility-based Perturbed Gradient Descent (UPGD) that performs gradient-based learning guided by utility-based information. The utility information is used as a gate, referred to as utility gating, for the gradients to prevent large updates to already useful weights, addressing forgetting. On the other hand, the utility information helps maintain plasticity by perturbing unuseful weights which become difficult to change through gradients (see Dohare et al. 2023a). Using shorthand, $\mathcal{L}=\mathcal{L}(\mathcal{W},Z)$, the update rule of UPGD is given by
@@
where $\xi\sim\mathcal{N}(0,\sigma^{2})$ is noise, $\alpha$ is the step-size parameter, and $\bar{U}_{l,i,j}\in[0,1]$ is a scaled utility. For important weights with utility $\bar{U}_{l,i,j}=1$, the weight remains unaltered even by gradient descent, whereas unimportant weights with $\bar{U}_{l,i,j}=0$ get updated by both perturbation and gradient descent.
@@ 4.2 UPGD Against Loss of Plasticity
We compare SGDW, PGD, S&P, which addresses loss of plasticity, Adam with weight decay (Loshchilov & Hutter 2019) known as AdamW, UPGD-W, and Non-protecting UPGD-W. We also introduce and compare against Streaming Elastic Weight Consolidation (S-EWC), Streaming Synaptic Intelligence (S-SI), and Streaming Memory-Aware Intelligence (S-MAS). These methods can be viewed as a natural extension of EWC (Kirkpatrick et al. 2017), SI (Zenke et al. 2017), and MAS (Aljundi et al. 2018), respectively, which are regularization-based methods for mitigating forgetting to the streaming learning setting. Finally, we introduce and compare against Streaming RWalk (S-RWalk). This can be seen as a natural extension of RWalk (Chaudhry et al. 2018), a method intended to address both issues, adapted for streaming learning. We write the update rule of the last four methods in streaming learning as $w_{l,i,j}\leftarrow w_{l,i,j}-\alpha\left[\nicefrac{{\partial\mathcal{L}}}{{\partial w_{l,i,j}}}+\kappa\Omega_{l,i,j}(w_{l,i,j}-\bar{w}_{l,i,j})\right]$, where $\kappa$ is the regularization factor, $\Omega_{l,i,j}$ is the weight importance, and $\bar{w}_{l,i,j}$ is the trace of weight $i,j$ at the $l$-th layer. The weight importance is estimated as a trace of squared gradients in S-EWC, whereas it is estimated as a trace of gradient magnitudes in S-MAS. S-RWalk is different from RWalk since it uses a trace of previous weights instead of the instantaneous previous weights. We omit the details of S-SI and RWalk weight importance estimation here and defer it to Appendix I.4. Lastly, since we compare against methods that use first-order information, we use global first-order utility traces in this and subsequent experiments for a fair comparison (however, see Appendix H for an experiment using second-order utilities).
@@ Appendix A Convergence Analysis for UPGD and Non-protecting UPGD
We further assume that $\mathcal{L}$ has bounded variance in the gradients $\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta},S)-\nabla f(\bm{\theta})\|^{2}]\leq\sigma_{g}^{2},\forall\bm{\theta}\in\mathbb{R}^{d}$. Note that the assumption of L-smoothness on the sample loss result in L-smooth expected loss too, which is given by $\|\nabla f(\bm{\theta}_{1})-\nabla f(\bm{\theta}_{2})\|\leq L\|\bm{\theta}_{1}-\bm{\theta}_{2}\|$. We assume that the perturbation noise in both UPGD versions has bounded variance $\mathbb{E}[\|\bm{\xi}\|^{2}]\leq\sigma^{2}_{n}$. For the simplicity of this proof, we use the true instantaneous weight utility, not an approximated one. We assume that the utility of any connection in the network is upper bounded by a number close to one $\bar{u}\to 1^{-}$.
@@ A.1 Non-protecting Utility-based Perturbed Gradient Descent
$\mathbb{E}[\|{\bm{g}}_{t}\|^{2}]$ $\leq\frac{\sigma_{g}^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2},$
@@
$\mathbb{E}_{t}[f(\bm{\theta}_{t+1})|\bm{\theta}_{t}]$ $\leq f(\bm{\theta}_{t})-\alpha\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2}\right)$
$=f(\bm{\theta}_{t})-\frac{2\alpha-L\alpha^{2}}{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}\right).$
@@
$\frac{2\alpha-L\alpha^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq f(\bm{\theta}_{1})-\mathbb{E}[f(\bm{\theta}_{T+1})]+\frac{L\alpha^{2}T}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}\right).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq 2\frac{f(\bm{\theta}_{1})-f(\bm{\theta}^{*})}{T(2\alpha-L\alpha^{2})}+\frac{L\alpha^{2}2(\sigma^{2}_{n}b_{t}+\sigma_{g}^{2})}{b_{t}(2\alpha-L\alpha^{2})}.$
@@ A.2 Utility-based Perturbed Gradient Descent
$\leq f(\bm{\theta}_{t})-\alpha\bar{\rho}\sum_{i=1}^{d}\nabla[f(\bm{\theta}_{t})]_{i}^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}+\|\nabla f(\bm{\theta}_{t})\|^{2}\right)$
$=f(\bm{\theta}_{t})-\left(\alpha\bar{\rho}-\frac{L\alpha^{2}}{2}\right)\|\nabla f(\bm{\theta}_{t})\|^{2}+\frac{L\alpha^{2}}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}\right).$
@@
$\frac{2\alpha\bar{\rho}-L\alpha^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq f(\bm{\theta}_{1})-\mathbb{E}[f(\bm{\theta}_{T+1})]+\frac{L\alpha^{2}T}{2}\left(\sigma^{2}_{n}+\frac{\sigma_{g}^{2}}{b_{t}}\right).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|^{2}$ $\leq 2\frac{f(\bm{\theta}_{1})-f(\bm{\theta}^{*})}{T(2\alpha\bar{\rho}-L\alpha^{2})}+\frac{L\alpha^{2}(\sigma^{2}_{n}b_{t}+\sigma_{g}^{2})}{b_{t}(2\alpha\bar{\rho}-L\alpha^{2})}.$
@@ Appendix B Utility Propagation
$f_{l,i,j}$ $\doteq\frac{\sigma_{a}^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}f_{l+1,k,i},$
$s_{l,i,j}$ $\doteq\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}2s_{l+1,k,i}\frac{\sigma_{a}^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma_{a}^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}.$
@@
$\frac{\partial\mathcal{L}}{\partial a_{l,i}}$ $=\sum_{k=1}^{|{\bm{a}}_{l+1}|}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}\frac{\partial a_{l+1,k}}{\partial h_{l,i}}\frac{\partial h_{l,i}}{\partial a_{l,i}}=\sigma_{a}^{\prime}(a_{l,i})\sum_{k=1}^{|{\bm{a}}_{l+1}|}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i},$ (9)
@@
$\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l,i}}}}$ $\doteq\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{[}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma_{a}^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma_{a}^{\prime\prime}(a_{l,i})\Bigg{]},$ (11)
@@
$f_{l,i,j}$ $=-\sigma_{a}^{\prime}(a_{l,i})h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\Bigg{)}$ (14)
$s_{l,i,j}$ $=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma_{a}^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma_{a}^{\prime\prime}(a_{l,i})\Bigg{)}$ (15)
@@
$f_{l,ij}$ $=-\sigma_{a}^{\prime}(a_{l,i})h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\Bigg{)}$
$=\frac{\sigma_{a}^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}-\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}h_{l,i}W_{l+1,k,i}\Bigg{)}$
$=\frac{\sigma_{a}^{\prime}(a_{l,i})}{h_{l,i}}h_{l-1,j}W_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}f_{l+1,k,i},$ (16)
$s_{l,i,j}$ $=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}W^{2}_{l+1,k,i}{\sigma_{a}^{\prime}}(a_{l,i})^{2}+\frac{\partial\mathcal{L}}{\partial a_{l+1,k}}W_{l+1,k,i}\sigma_{a}^{\prime\prime}(a_{l,i})\Bigg{)}$
$=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}\widehat{\frac{\partial^{2}\mathcal{L}}{\partial{a^{2}_{l+1,k}}}}h^{2}_{l,i}W^{2}_{l+1,k,i}\frac{\sigma_{a}^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma_{a}^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}$
$=\frac{1}{2}h^{2}_{l-1,j}W^{2}_{l,i,j}\sum_{k=1}^{|{\bm{a}}_{l+1}|}\Bigg{(}2s_{l+1,k,i}\frac{\sigma_{a}^{\prime}(a_{l,i})^{2}}{h^{2}_{l,i}}-\frac{\sigma_{a}^{\prime\prime}(a_{l,i})}{h_{l,i}}f_{l+1,k,i}\Bigg{)}.$ (17)